# Learning Theory for Kernel Bilevel Optimization

### v1

@@ 2.3 Implicit Differentiation in an RKHS
A stationarity measure in (KBO) is the gradient $\nabla\mathcal{F}(\omega)$ of the value function $\mathcal{F}$. Nonetheless, evaluating the gradient requires computing the Jacobian $\partial_{\omega}h^{\star}_{\omega}$, which can be viewed as a linear operator from $\mathcal{H}$ to $\mathbb{R}^{d}$. Indeed $h^{\star}_{\omega}$ depends implicitly on $\omega$. A key ingredient for computing the Jacobian $\partial_{\omega}h^{\star}_{\omega}$ is the implicit function theorem (Ioffe & Tihomirov, 1979) which guarantees differentiability of the implicit function $\omega\mapsto h_{\omega}^{\star}$ and allows characterizing $\partial_{\omega}h^{\star}_{\omega}$ as the unique solution of a linear system of the form:
@@
where $\partial_{h}^{2}L_{in}(\omega,h^{\star}_{\omega})$ is a linear operator from $\mathcal{H}$ to itself representing the partial Hessian of $L_{in}$ w.r.t. $h$, while $\partial_{\omega,h}^{2}L_{in}(\omega,h^{\star}_{\omega})$ is an operator from $\mathcal{H}$ to $\mathbb{R}^{d}$ representing the cross derivatives of $L_{in}$ w.r.t. to $\omega$ and $h$. Applying such result requires $h\mapsto L_{in}(\omega,h)$ to be Fréchet differentiable with gradient map $(\omega,h)\mapsto\partial_{h}L_{in}(\omega,h)$ jointly Fréchet differentiable and invertible Hessian operator. All these properties are satisfied in our setting under Assumptions (A), (B), (D) and (C) as shown in Propositions A.1, A.2 and A.3 of Section A.1. Furthermore, when the outer objective $L_{out}$ is Fréchet differentiable, which is our case under our assumptions (Proposition A.1 of Section A.1), then by composition with $\omega\mapsto(\omega,h_{\omega}^{\star})$, the map $\omega\mapsto\mathcal{F}(\omega)$ must also be differentiable with gradient obtained using the chain rule:
$\nabla\mathcal{F}(\omega)=\partial_{\omega}L_{out}(\omega,h_{\omega}^{\star})-\partial_{\omega}h_{\omega}^{\star}\partial_{h}L_{out}(\omega,h_{\omega}^{\star}).$
The above expression for the gradient is intractable as it involves abstract operators. In Proposition 2.1 below, we derive an explicit expression for $\nabla\mathcal{F}(\omega)$ which exploits the particular structure of the objectives $L_{in}$ and $L_{out}$ as expectations of point-wise losses.
@@
Under Assumptions (B), (D), (A) and (C), $\mathcal{F}$ is differentiable on $\mathbb{R}^{d}$, with gradient $\nabla\mathcal{F}(\omega)$, for any $\omega\in\mathbb{R}^{d}$, given by:
$\nabla\mathcal{F}(\omega)=$ $\mathbb{E}_{\mathbb{Q}}\left[\partial_{\omega}\ell_{out}(\omega,h^{\star}_{\omega}(x),y)\right]$ (2)
$+\mathbb{E}_{\mathbb{P}}\left[\partial_{\omega,v}^{2}\ell_{in}(\omega,h^{\star}_{\omega}(x),y)a^{\star}_{\omega}(x)\right],$
where the adjoint function $a^{\star}_{\omega}\in\mathcal{H}$ is the unique minimizer of a strongly-convex quadratic objective $a\mapsto L_{adj}(\omega,a)$ defined on $\mathcal{H}$ as:
$\begin{split}L_{adj}(\omega,&a)\coloneqq\frac{1}{2}\mathbb{E}_{\mathbb{P}}{\left[\partial_{v}^{2}\ell_{in}{\left(\omega,h_{\omega}^{\star}(x),y\right)}a(x)^{2}\right]}\\ &+\mathbb{E}_{\mathbb{Q}}{\left[\partial_{v}\ell_{out}{\left(\omega,h_{\omega}^{\star}(x),y\right)}a(x)\right]}+\frac{\lambda}{2}{\left\lVert a\right\rVert}_{\mathcal{H}}^{2},\end{split}$ (3)
@@
Proposition 2.1 is proved in Section A.1 and relies essentially on proving Bochner’s integrability (Diestel & Uhl, 1977, Definition 1, Chapter 2) of some suitable operators on $\mathcal{H}$, and then applying Lebesgue’s dominated convergence theorem for Bochner’s integral (Diestel & Uhl, 1977, Theorem 3, Chapter 2) to exchange derivatives and expectations. The expression in Proposition 2.1 provides a natural way for approximating $\nabla\mathcal{F}(\omega)$ by estimating all expectations using finite sample averages, as we further discuss in Section 3.
3 Finite Sample Approximation of (KBO)
In this section, we consider an approximation to (KBO) problem when only a finite number of i.i.d. samples $(x_{i},y_{i})_{1\leq i\leq n}$ and $(\tilde{x}_{j},\tilde{y}_{j})_{1\leq j\leq m}$ from $\mathbb{P}$ and $\mathbb{Q}$ are available. This setting is ubiquitous in machine learning as it allows finding tractable approximate solutions to the original problem. As we are interested in approximately solving (KBO) using gradient methods, our focus here is to derive estimators for both the value function $\mathcal{F}(\omega)$ and its gradient $\nabla\mathcal{F}(\omega)$ for which the generalization properties will later be studied in Section 4.
[…]
@@ 3.1 Value Function: Plug-in Estimator and its Gradient
$\begin{split}\widehat{L}_{out}{\left(\omega,h\right)}&\coloneqq\frac{1}{m}\sum_{j=1}^{m}\ell_{out}(\omega,h(\tilde{x}_{j}),\tilde{y}_{j})\\ \widehat{L}_{in}{\left(\omega,h\right)}&\coloneqq\frac{1}{n}\sum_{i=1}^{n}\ell_{in}(\omega,h(x_{i}),y_{i})+\frac{\lambda}{2}\|h\|_{\mathcal{H}}^{2}.\end{split}$
A plug-in estimator $\omega\mapsto\widehat{\mathcal{F}}(\omega)$ is then obtained by first finding a solution $\hat{h}_{\omega}$ minimizing $h\mapsto\widehat{L}_{in}(\omega,h)$ that is meant to approximate the optimal inner solution $h_{\omega}^{\star}$, and subsequently plugging it into $\widehat{L}_{out}$. This procedure results in the following empirical version of (KBO):
$\begin{split}&\min_{\omega\in\mathcal{C}}\widehat{\mathcal{F}}(\omega)\coloneqq\widehat{L}_{out}{\left(\omega,\hat{h}_{\omega}\right)}\\ &\text{s.t.}\quad\hat{h}_{\omega}=\operatorname*{arg\,min}_{h\in\mathcal{H}}\widehat{L}_{in}{\left(\omega,h\right)}.\end{split}$
The inner problem still requires optimizing over a, potentially infinite-dimensional, RKHS. However, its finite sum structure allows equivalently expressing it as a finite-dimensional bilevel optimization, by application of the so called representer theorem (Schölkopf et al., 2001):
@@
[…]
In the above equation, $\operatorname{\mathbf{K}}\in\mathbb{R}^{n\times n}$ and $\operatorname{\mathbf{\overline{K}}}\in\mathbb{R}^{m\times n}$ are matrices containing the pairwise kernel similarities between the data points, i.e., $\operatorname{\mathbf{K}}_{ij}\coloneqq K(x_{i},x_{j})$ and $\operatorname{\mathbf{\overline{K}}}_{ij}\coloneqq K(\tilde{x}_{i},x_{j})$, while $\gamma$ is a parameter vector in $\mathbb{R}^{n}$ representing the inner-level variables. The optimal solution $\hat{\gamma}_{\omega}$ enables recovering the prediction function $\hat{h}_{\omega}$ by linearly combining kernel evaluations at inner-level samples, i.e., $\hat{h}_{\omega}=\sum_{i=1}^{n}(\hat{\gamma}_{\omega})_{i}K(x_{i},\cdot)$.
@@
The formulation in ($\widehat{\text{KBO}}$) allows deriving an expression for the gradient ^$\nabla\widehat{\mathcal{F}}(\omega)$ in terms of the Jacobian $\partial_{\omega}\hat{\gamma}_{\omega}$ by direct application of the chain rule. Unlike the Jacobian $\partial_{\omega}h_{\omega}^{\star}$ which requires solving an infinite-dimensional linear system given by Equation 1, the Jacobian of $\hat{\gamma}_{\omega}$ can be obtained as a solution of a finite-dimensional linear system by application of the implicit function theorem (see Proposition B.1 of Appendix B). Hence, ($\widehat{\text{KBO}}$) falls into a class of optimization problems for which a rich body of literature have proposed practical and scalable algorithms, leveraging the expression of ^$\nabla\widehat{\mathcal{F}}(\omega)$ (Ji et al., 2021; Arbel & Mairal, 2022a; Dagréou et al., 2022). Consequently, solving ($\widehat{\text{KBO}}$) provides a practical way for approximating the solution to the original population problem (KBO) as proposed in several prior works on bilevel optimization involving kernel methods (Keerthi et al., 2006; Kunapuli et al., 2008).
@@
Despite its practical advantage, the above approach yields algorithms that are not directly amenable to a statistical analysis. The key challenge is to be able to control the approximation error between the true gradient $\nabla\mathcal{F}(\omega)$ and its approximation ^$\nabla\widehat{\mathcal{F}}(\omega)$ as the sample sizes $n$ and $m$ increase. Existing statistical analysis for bilevel optimization, such as (Bao et al., 2021; Zhang et al., 2024), considered objectives in the form of expectations/finite sums of point-wise losses, as we do here. However, they require both inner-level and outer-level parameters to belong to spaces of fixed dimensions that is independent of the sample sizes $n$ and $m$. That is because these parameters are expected to converge towards some fixed vectors as $n,m\rightarrow+\infty$. Unfortunately, these results are not applicable in our case since the inner-level parameter $\gamma$ has a dimension that increases with the sample size $n$ ($\gamma\in\mathbb{R}^{n}$) and is not expected to converge towards any well-defined object. Next, we provide an equivalent expression for ^$\nabla\widehat{\mathcal{F}}(\omega)$ that will be crucial in our statistical analysis in Section 4.
@@ 3.2 Plug-in Estimator of the Total Gradient
We consider now an, a priori, different approach for approximating the total gradient $\nabla\mathcal{F}(\omega)$ based on direct plug-in estimation from Equation 2 and show that it recovers the previously introduced estimator ^$\nabla\widehat{\mathcal{F}}(\omega)$. Such approach consists in replacing all expectations in Equation 2 by empirical averages, then replacing $h^{\star}_{\omega}$ and $a_{\omega}^{\star}$ by finite sample estimates $\hat{h}_{\omega}$ and $\hat{a}_{\omega}$:
^$\widehat{\nabla\mathcal{F}}(\omega)=$ $\frac{1}{m}\sum_{j=1}^{m}\partial_{\omega}\ell_{out}(\omega,\hat{h}_{\omega}(\tilde{x}_{j}),\tilde{y}_{j})$ (4)
$+\frac{1}{n}\sum_{i=1}^{n}\partial_{\omega,v}^{2}\ell_{in}(\omega,\hat{h}_{\omega}(x_{i}),y_{i})\hat{a}_{\omega}(x_{i}).$
Just as in Section 3.1, $h^{\star}_{\omega}$ can be estimated by $\hat{h}_{\omega}$, the minimizer of the empirical objective $h\mapsto\widehat{L}_{in}(\omega,h)$. Similarly, $a^{\star}_{\omega}$ can be estimated by minimizing an empirical version $a\mapsto\widehat{L}_{adj}(\omega,a)$ of the adjoint objective $L_{adj}$ given in Equation 3:
$\widehat{L}_{adj}(\omega,a)=\frac{1}{2n}\sum_{i=1}^{n}\partial_{v}^{2}\ell_{in}(\omega,\hat{h}_{\omega}(x_{i}),y_{i})a^{2}(x_{i})$ (5)
$+\frac{1}{m}\sum_{j=1}^{m}\partial_{v}\ell_{out}(\omega,\hat{h}_{\omega}(\tilde{x}_{j}),\tilde{y}_{j})a(\tilde{x}_{j})+\frac{\lambda}{2}\|a\|_{\mathcal{H}}^{2}.$
Both functions $\hat{h}_{\omega}$ and $\hat{a}_{\omega}$ can be expressed as linear combinations of kernel evaluations with some given parameters that increase with the sample size $n$, (see Proposition B.1 for $\hat{h}_{\omega}$ and Lemma B.2 for $\hat{a}_{\omega}$, both in Appendix B). However, these parameters are never required for expressing the plug-in estimator ^$\widehat{\nabla\mathcal{F}}(\omega)$ in Equation 4, since only the function values of $\hat{h}_{\omega}$ and $\hat{a}_{\omega}$ are needed for computing it. This property is precisely what makes ^$\widehat{\nabla\mathcal{F}}(\omega)$ suitable for a statistical analysis, since its estimation error depends on the errors of $\hat{h}_{\omega}$ and $\hat{a}_{\omega}$ which always belong to the same space $\mathcal{H}$, regardless of the sample size, and are expected to approach their population counter-parts. This is unlike the expression of ^$\nabla\widehat{\mathcal{F}}(\omega)$ obtained by implicit differentiation and which suggests controlling the behavior of the vector $\hat{\gamma}_{\omega}$.
The next proposition demonstrates that, surprisingly, both estimators ^$\nabla\widehat{\mathcal{F}}(\omega)$ and ^$\widehat{\nabla\mathcal{F}}(\omega)$ are precisely equal.
@@
Proposition 3.1 is proved in Appendix B and relies on an application of the representer theorem (Schölkopf et al., 2001) to provide explicit expressions for both estimators in terms of $\gamma_{\omega}$, kernel matrices $\operatorname{\mathbf{K}}$ and $\operatorname{\mathbf{\overline{K}}}$ and partial derivatives of the point-wise objectives $\ell_{in}$ and $\ell_{out}$. Both expressions are then shown to be equal using optimality conditions on the parameters defining $\hat{a}_{\omega}$. The result in Proposition 3.1 precisely says that the operations of differentiation and plug-in estimation commute in the case of (KBO). Such a commutativity property does not necessarily hold anymore if one considers spaces other than an RKHS as discussed in (Petrulionytė et al., 2024, Appendix F). Next, we leverage the expression of the plug-in estimator ^$\widehat{\nabla\mathcal{F}}(\omega)$ to provide generalization bounds.

### v3

@@ 2.3 Implicit differentiation in an RKHS
A stationarity measure in (KBO) is the gradient $\nabla\mathcal{F}(\omega)$ of the value function $\mathcal{F}$. Computing $\nabla\mathcal{F}(\omega)$, however, is challenging and will be addressed in this section. At a high level, our approach proceeds in two steps. First, we derive an abstract, a priori intractable, expression for $\nabla\mathcal{F}(\omega)$ using implicit differentiation in an RKHS, which is the main source of difficulty. Then, we leverage the structure of our problem to reformulate the gradient in Proposition˜2.2 using the solution of a regression problem in the RKHS (the adjoint problem). This more concrete formulation can be approximated with finite samples and will later serve as the foundation of our statistical analysis. Formally, evaluating the gradient requires computing the Jacobian $\partial_{\omega}h^{\star}_{\omega}$, which can be viewed as a linear operator from $\mathcal{H}$ to $\mathbb{R}^{d}$. Indeed, $h^{\star}_{\omega}$ depends implicitly on $\omega$. A key ingredient for computing $\partial_{\omega}h^{\star}_{\omega}$ is the implicit function theorem [ioffe1979theory], which guarantees the differentiability of the implicit function $\omega\mapsto h_{\omega}^{\star}$ and allows characterizing $\partial_{\omega}h^{\star}_{\omega}$ as the unique solution of a linear system of the form:
@@
where $\partial_{h}^{2}L_{in}(\omega,h^{\star}_{\omega})$ is an operator from $\mathcal{H}$ to itself representing the Hessian of $L_{in}$ w.r.t. $h$, while $\partial_{\omega,h}^{2}L_{in}(\omega,h^{\star}_{\omega})$ is an operator from $\mathcal{H}$ to $\mathbb{R}^{d}$ representing the cross derivatives of $L_{in}$ w.r.t. to $\omega$ and $h$. Applying such result requires $h\mapsto L_{in}(\omega,h)$ to be Fréchet differentiable with invertible Hessian operator and jointly Fréchet differentiable gradient map $(\omega,h)\mapsto\partial_{h}L_{in}(\omega,h)$. All these properties are satisfied in our setting under Assumptions˜(A), (B), (C), (E) and (D) as shown in Propositions˜B.1, B.2 and B.3 of Section˜B.1. Furthermore, when $L_{out}$ is Fréchet differentiable, which is our case under Assumptions˜(A), (B), (C) and (D) (Proposition˜B.1 of Section˜B.1), then by composition with $\omega\mapsto(\omega,h_{\omega}^{\star})$, the map $\omega\mapsto\mathcal{F}(\omega)$ must also be differentiable with gradient obtained using the chain rule:
$\nabla\mathcal{F}(\omega)=\partial_{\omega}L_{out}(\omega,h_{\omega}^{\star})+\partial_{\omega}h_{\omega}^{\star}\partial_{h}L_{out}(\omega,h_{\omega}^{\star}).$
The above expression for the gradient is intractable as it involves abstract operators, namely the derivatives $\partial_{h}$, $\partial_{h}^{2}$, and $\partial_{\omega,h}^{2}$, the last two of which arise when replacing $\partial_{\omega}h^{\star}_{\omega}$ by its expression in Equation˜1. In Proposition˜2.2 below, we derive an explicit expression for $\nabla\mathcal{F}(\omega)$ which exploits the particular structure of the objectives $L_{in}$ and $L_{out}$ as expectations of point-wise losses.
@@
Under Assumptions˜(A), (C), (E), (B) and (D), $\mathcal{F}$ is differentiable on $\mathbb{R}^{d}$, with gradient $\nabla\mathcal{F}(\omega)$, for any $\omega\in\mathbb{R}^{d}$, given by:
$\nabla\mathcal{F}(\omega)=\mathbb{E}_{\mathbb{Q}}\left[\partial_{\omega}\ell_{out}(\omega,h^{\star}_{\omega}(x),y)\right]+\mathbb{E}_{\mathbb{P}}\left[\partial_{\omega,v}^{2}\ell_{in}(\omega,h^{\star}_{\omega}(x),y)a^{\star}_{\omega}(x)\right],$ (2)
where the adjoint function $a^{\star}_{\omega}\in\mathcal{H}$ is the unique minimizer of a strongly convex quadratic objective $a\mapsto L_{adj}(\omega,a)$ defined on $\mathcal{H}$ as:
$L_{adj}(\omega,$ $a)\coloneqq\frac{1}{2}\mathbb{E}_{\mathbb{P}}\left[\partial_{v}^{2}\ell_{in}\left(\omega,h_{\omega}^{\star}(x),y\right)a^{2}(x)\right]+\mathbb{E}_{\mathbb{Q}}\left[\partial_{v}\ell_{out}\left(\omega,h_{\omega}^{\star}(x),y\right)a(x)\right]+\frac{\lambda}{2}\left\lVert a\right\rVert_{\mathcal{H}}^{2},$ (3)
@@
Proposition˜2.2 is proved in Appendix˜B and relies essentially on proving Bochner’s integrability [diestel1977vector, Definition 1, Chapter 2] of some suitable operators on $\mathcal{H}$, and then applying Lebesgue’s dominated convergence theorem for Bochner’s integral [diestel1977vector, Theorem 3, Chapter 2] to interchange derivatives and expectations. The expression in Proposition˜2.2 provides a natural way for approximating $\nabla\mathcal{F}(\omega)$ by estimating all expectations using finite-sample averages, as we further discuss in Section˜3.
3 Finite-sample approximation of (KBO)
@@
In this section, we consider an approximation of (KBO) when only a finite number of i.i.d. samples $(x_{i},y_{i})_{1\leq i\leq n}$ and $(\tilde{x}_{j},\tilde{y}_{j})_{1\leq j\leq m}$ from $\mathbb{P}$ and $\mathbb{Q}$ are available. This setting is ubiquitous in machine learning as it allows finding tractable approximate solutions to the original problem. As we are interested in approximately solving (KBO) using gradient methods, our focus here is to derive estimators for both the value function $\mathcal{F}(\omega)$ and its gradient $\nabla\mathcal{F}(\omega)$, whose generalization properties will be studied in Section˜4.
[…]
@@ 3.1 Value function: plug-in estimator and its gradient
$\widehat{L}_{out}\left(\omega,h\right)=\frac{1}{m}\sum_{j=1}^{m}\ell_{out}(\omega,h(\tilde{x}_{j}),\tilde{y}_{j}),\quad\widehat{L}_{in}\left(\omega,h\right)=\frac{1}{n}\sum_{i=1}^{n}\ell_{in}(\omega,h(x_{i}),y_{i})+\frac{\lambda}{2}\|h\|_{\mathcal{H}}^{2}.$
A plug-in estimator $\omega\mapsto\widehat{\mathcal{F}}(\omega)$ is then obtained by first finding a solution $\hat{h}_{\omega}$ minimizing $h\mapsto\widehat{L}_{in}(\omega,h)$, that is meant to approximate the optimal inner solution $h_{\omega}^{\star}$, and subsequently plugging it into $\widehat{L}_{out}$. This procedure results in the following empirical version of (KBO):
$\min_{\omega\in\mathcal{C}}\widehat{\mathcal{F}}(\omega)\coloneqq\widehat{L}_{out}(\omega,\hat{h}_{\omega})\quad\text{s.t.}\quad\hat{h}_{\omega}=\operatorname*{arg\,min}_{h\in\mathcal{H}}\widehat{L}_{in}\left(\omega,h\right).$
The inner problem still requires optimizing over a, potentially infinite-dimensional, RKHS. However, its finite-sum structure allows equivalently expressing it as a finite-dimensional bilevel optimization, by application of the so-called representer theorem [scholkopf2001generalized]:
@@
$\text{s.t.}\quad\hat{\operatorname{\boldsymbol{\gamma}}}_{\omega}$ $=\operatorname*{arg\,min}_{\operatorname{\boldsymbol{\gamma}}\in\mathbb{R}^{n}}\frac{1}{n}\sum_{i=1}^{n}\ell_{in}(\omega,\left(\operatorname{\mathbf{K}}\operatorname{\boldsymbol{\gamma}}\right)_{i},y_{i})+\frac{\lambda}{2}\operatorname{\boldsymbol{\gamma}}^{\top}\operatorname{\mathbf{K}}\operatorname{\boldsymbol{\gamma}}.$
@@
Here, $\operatorname{\mathbf{K}}\in\mathbb{R}^{n\times n}$ and $\operatorname{\mathbf{\overline{K}}}\in\mathbb{R}^{m\times n}$ are matrices containing the pairwise kernel similarities between the data points, i.e., $\operatorname{\mathbf{K}}_{ij}\coloneqq K(x_{i},x_{j})$ and $\operatorname{\mathbf{\overline{K}}}_{ij}\coloneqq K(\tilde{x}_{i},x_{j})$, while $\operatorname{\boldsymbol{\gamma}}$ is a parameter vector in $\mathbb{R}^{n}$ representing the inner-level variables. The optimal solution $\hat{\operatorname{\boldsymbol{\gamma}}}_{\omega}$ enables recovering the prediction function $\hat{h}_{\omega}$ by linearly combining kernel evaluations at inner-level samples, i.e., $\hat{h}_{\omega}=\sum_{i=1}^{n}(\hat{\operatorname{\boldsymbol{\gamma}}}_{\omega})_{i}K(x_{i},\cdot)$. The formulation in ($\widehat{\text{KBO}}$) enables deriving an expression for the gradient $\nabla\widehat{\mathcal{F}}(\omega)$ in terms of the Jacobian $\partial_{\omega}\hat{\operatorname{\boldsymbol{\gamma}}}_{\omega}$ by direct application of the chain rule. Unlike $\partial_{\omega}h_{\omega}^{\star}$, which requires solving the infinite-dimensional linear system in Equation˜1, $\partial_{\omega}\hat{\operatorname{\boldsymbol{\gamma}}}_{\omega}$ can be obtained by solving a finite-dimensional linear system using the implicit function theorem (see Proposition˜C.1 of Appendix˜C). Hence, ($\widehat{\text{KBO}}$) falls into a class of optimization problems for which a rich body of literature has proposed practical and scalable algorithms, leveraging the expression of $\nabla\widehat{\mathcal{F}}(\omega)$ [Ji:2021a, Arbel:2021a, Dagreou2022SABA]. Solving ($\widehat{\text{KBO}}$) thus provides a practical way to approximate the solution of the original population problem (KBO), as proposed by several prior works on bilevel optimization involving kernel methods [keerthi2006efficient, kunapuli2008classification].
@@
Non-applicability of existing results. Despite its practical advantages, the above approach yields algorithms that are not directly amenable to a statistical analysis. The key challenge is to be able to control the approximation error between the true gradient $\nabla\mathcal{F}(\omega)$ and its approximation $\nabla\widehat{\mathcal{F}}(\omega)$ as the sample sizes $n$ and $m$ increase. Existing statistical analyses for bilevel optimization, such as [bao2021stability, zhang2024fine], consider objectives that are expectations or finite sums of point-wise losses, as we do here. While these results can be applied to our setting for each fixed $n$, they do not capture the generalization behavior as $n$ grows. In particular, they require both the inner- and outer-level parameters to lie in spaces of fixed dimensions, that are independent of $n$ and $m$. That is because these parameters are expected to converge to some fixed vectors as $n,m\rightarrow+\infty$. In contrast, in our setting, the inner-level parameter $\operatorname{\boldsymbol{\gamma}}$ lies in $\mathbb{R}^{n}$, so its dimension grows with $n$ and is not expected to converge to any well-defined object. Existing non-kernel generalization bounds are discussed in Appendix˜A.
Relation to instrumental variable regression. Our formulation is related to instrumental variable regression, which is a special case, but differs in that we study regularized inner problems with a fixed $\lambda$, independent of the sample sizes. In contrast, the instrumental variable regression literature typically considers un-regularized population problems ($\lambda=0$), for which a closed-form expression of the inner minimizer is available [pmlr-v70-hartford17a, singh2019kernel, xu2021learning, li2024regularized]. Our contribution lies in an orthogonal direction: we handle more general inner objectives, for which even the analysis of regularized problems raises new obstacles that had not been tackled before. Moreover, some prior works have provided convergence rates in the instrumental variable regression setting [ai2003efficient, ai2007estimation, chen2012estimation, chen2015sieve]. Yet, these studies focus on asymptotic results with sieve estimators, meanwhile we leverage the RKHS structure to provide finite-sample bounds. More recently, meunier2024nonparametric established minimax optimal rates under source assumptions by exploiting bounds for vector-valued kernel ridge regression [li2024optimal] via a spectral filtering technique [meunier2024optimal]. However, this approach is not applicable to our case, as it requires the losses to be quadratic in $\omega$, as further discussed in Appendix˜A.
Next, we provide an equivalent expression for $\nabla\widehat{\mathcal{F}}(\omega)$, essential for our analysis in Section˜4.
@@ 3.2 Plug-in estimator of the total gradient
We now consider an a priori different approach for approximating the total gradient $\nabla\mathcal{F}(\omega)$ based on direct plug-in estimation from Equation˜2, and show that it recovers the previously introduced estimator $\nabla\widehat{\mathcal{F}}(\omega)$. Such approach consists in replacing all expectations in Equation˜2 by empirical averages, then replacing $h^{\star}_{\omega}$ and $a_{\omega}^{\star}$ by their finite-sample estimates $\hat{h}_{\omega}$ and $\hat{a}_{\omega}$. This yields the following estimator of the total gradient:
$\widehat{\nabla\mathcal{F}}(\omega)=\frac{1}{m}\sum_{j=1}^{m}\partial_{\omega}\ell_{out}(\omega,\hat{h}_{\omega}(\tilde{x}_{j}),\tilde{y}_{j})+\frac{1}{n}\sum_{i=1}^{n}\partial_{\omega,v}^{2}\ell_{in}(\omega,\hat{h}_{\omega}(x_{i}),y_{i})\hat{a}_{\omega}(x_{i}).$ (4)
Just as in Section˜3.1, $h^{\star}_{\omega}$ can be estimated by $\hat{h}_{\omega}$, the minimizer of the empirical objective $h\mapsto\widehat{L}_{in}(\omega,h)$. Similarly, $a^{\star}_{\omega}$ can be approximated by $\hat{a}_{\omega}$, the minimizer of $a\mapsto\widehat{L}_{adj}(\omega,a)$ defined as:
$\widehat{L}_{adj}(\omega,a)=\frac{1}{2n}\sum_{i=1}^{n}\partial_{v}^{2}\ell_{in}(\omega,\hat{h}_{\omega}(x_{i}),y_{i})a^{2}(x_{i})+\frac{1}{m}\sum_{j=1}^{m}\partial_{v}\ell_{out}(\omega,\hat{h}_{\omega}(\tilde{x}_{j}),\tilde{y}_{j})a(\tilde{x}_{j})+\frac{\lambda}{2}\|a\|_{\mathcal{H}}^{2},$ (5)
which serves as the empirical counterpart of the adjoint objective $L_{adj}$ given in Equation˜3. Both functions $\hat{h}_{\omega}$ and $\hat{a}_{\omega}$ can be expressed as linear combinations of kernel evaluations with some given parameter vectors whose dimensions increase with the sample size $n$ (see Proposition˜C.1 for $\hat{h}_{\omega}$ and Lemma˜C.2 for $\hat{a}_{\omega}$, both in Appendix˜C). However, these parameters are not required to compute the plug-in estimator $\widehat{\nabla\mathcal{F}}(\omega)$ in Equation˜4, since only the function values of $\hat{h}_{\omega}$ and $\hat{a}_{\omega}$ are needed. This property is precisely what makes $\widehat{\nabla\mathcal{F}}(\omega)$ suitable for a statistical analysis. Indeed, its estimation error depends on the approximation errors of $\hat{h}_{\omega}$ and $\hat{a}_{\omega}$, which always belong to the same space $\mathcal{H}$ regardless of the sample size, and are expected to approach their population counterparts. This contrasts with $\nabla\widehat{\mathcal{F}}(\omega)$ obtained by implicit differentiation, whose analysis would need controlling the behavior of the vector $\hat{\operatorname{\boldsymbol{\gamma}}}_{\omega}$ that resides in a growing-dimensional space as $n\to+\infty$.
The next proposition establishes a link between practical applications and theoretical analysis by demonstrating that, surprisingly, both estimators $\nabla\widehat{\mathcal{F}}(\omega)$ and $\widehat{\nabla\mathcal{F}}(\omega)$ are precisely equal.
@@
Proposition˜3.1 is proved in Appendix˜C and relies on an application of the representer theorem [scholkopf2001generalized] to provide explicit expressions for both estimators in terms of $\hat{\operatorname{\boldsymbol{\gamma}}}_{\omega}$, kernel matrices $\operatorname{\mathbf{K}}$ and $\operatorname{\mathbf{\overline{K}}}$ and partial derivatives of the point-wise objectives $\ell_{in}$ and $\ell_{out}$. Both expressions are then shown to be equal using optimality conditions on the parameters defining $\hat{a}_{\omega}$. The result in Proposition˜3.1 precisely says that the operations of differentiation and plug-in estimation commute in the case of (KBO). Such a commutativity property does not necessarily hold anymore if one considers spaces other than an RKHS, such as $L_{2}$ [petrulionyte2024functional, Appendix F]. The main difficulty arises from the argmin constraint and the use of implicit differentiation, which may introduce non-linear dependencies between inner- and outer-level variables, making the exchange of differentiation and discretization nontrivial. Next, we leverage the expression of the plug-in estimator $\widehat{\nabla\mathcal{F}}(\omega)$ to provide generalization bounds.